OpenStudy (anonymous):

Step 1: |x - 5| + 2 = 5 Step 2: |x - 5| = 5 - 2 Step 3: |x - 5| = 3

OpenStudy (anonymous):

help anyone?

OpenStudy (amistre64):

what absolute values get us to +3 ?

OpenStudy (anonymous):

one bigger than five?

OpenStudy (anonymous):

well a number bigger than 5?

OpenStudy (amistre64):

|3| = 3 and |-3| = 3 would you agree?

OpenStudy (anonymous):

oh ya

OpenStudy (amistre64):

then the next step is to split the innards: x-5 = 3 or x-5 = -3

OpenStudy (anonymous):

yes I know that but idk what the next step is to get the answer

OpenStudy (amistre64):

add 5 of course .... algebra doesnt change just becuase you split something into 2 possibilities

OpenStudy (anonymous):

ohhh ok so it would be x+5=3 OR x+5=-3?

OpenStudy (amistre64):

not quite, but prolly a mistype x-5 = 3 or x-5 = -3 x = 3+5 or x = -3+5

OpenStudy (anonymous):

but why x before the = sign now?

OpenStudy (amistre64):

becuase -5+5 = 0

OpenStudy (amistre64):

your previous steps assumes that you know how to work the sides of an equation

OpenStudy (anonymous):

I don't understand?.......

OpenStudy (amistre64):

x-5 = 3, solve for x x-5 = -3, solve for x

OpenStudy (amistre64):

adding 5 to both sides gives us what?

OpenStudy (anonymous):

7 or 2

OpenStudy (amistre64):

8 or 2, but yeah

OpenStudy (amistre64):

Step 1: |x - 5| + 2 = 5 Step 2: |x - 5| = 5 - 2 Step 3: |x - 5| = 3 step4: split the innards x-5 = 3 or x-5 = -3 step 5: x-5+5 = 3+5 or x-5+5 = -3+5